Using the following chemical equation how many grams of carbon dioxide can be produced by burning 18.49 g of butane?

C4H10 + 5O2 -> 3CO2 + 4H2O


1
Expert's answer
2021-02-16T07:41:42-0500

According to the equation, the mass of carbon dioxide is:

m(CO2) = m(C4H10) × 3 × Mr(CO2) / Mr(C4H10) = 18.49 g × 3 × 44 g/mol / 58 g/mol = 42.08 g


Answer: 42.08 g